OpenStudy (firejay5):

Solve each equation. Check your solution.: Please be clear on your answer and explain how you got it. 3/4 - 1/2X = 5/8

OpenStudy (lalaly):

is it\[\frac{3}{4}-\frac{1}{2}x=\frac{5}{8}\]?

OpenStudy (firejay5):

yes

OpenStudy (lalaly):

subtract 3/4 from both sides\[-\frac{1}{2}x=\frac{5}{8}-\frac{3}{4}\]\[-\frac{1}{2}x=-\frac{1}{8}\]multiply both sides by -2\[x=\frac{2}{8}=\frac{1}{4}\]

OpenStudy (firejay5):

nevermind